Test for a positive correlation using the sample correlation r=0.33 and the sample size n=30.

State the null and alternative hypotheses.

  

  

Find the test statistic.

Round your answer to two decimal places.

t=


Find the p-value.

Round your answer to three decimal places.

The p-value is  .


What is the conclusion, using a 5% significance level?

                                                                      Do not rejectReject H0.


Is the correlation significant?

                                                                      NoYes
Answer #1

Hypothesis: Vs  

Here , use t-test.

df=degrees of freedom=n-1=30-1=29

The test statistic is ,

The p-value is ,

p-value= ; The Excel function is , =TDIST(1.85,28,1)

Decision : Here , p-value=0.037< 0.05

Therefore ,reject Ho.

Conclusion : Yes , the correlation is significant.